OpenStudy (anonymous):

2/v=9/v-1/2 The solution is 14? Right?

OpenStudy (anonymous):

can u write ur question clearly please i cannot get ur question!!

OpenStudy (anonymous):

Yeap

OpenStudy (anonymous):

My problem wants me to find the solution. \[\left(\begin{matrix}2 \\ v\end{matrix}\right)=\left(\begin{matrix}9 \\ v\end{matrix}\right)-\left(\begin{matrix}1 \\ 2\end{matrix}\right)\]

OpenStudy (anonymous):

Thank you roshan2004!!

OpenStudy (anonymous):

Yes.